# Sparse Voxels Rasterization: Real-time High-fidelity Radiance Field Rendering

### v1

@@ Abstract
We propose an efficient radiance field rendering algorithm that incorporates a rasterization process on sparse voxels without neural networks or 3D Gaussians. There are two key contributions coupled with the proposed system. The first is to render sparse voxels in the correct depth order along pixel rays by using dynamic Morton ordering. This avoids the well-known popping artifact found in Gaussian splatting. Second, we adaptively fit sparse voxels to different levels of detail within scenes, faithfully reproducing scene details while achieving high rendering frame rates. Our method improves the previous neural-free voxel grid representation by over 4db PSNR and more than 10x rendering FPS speedup, achieving state-of-the-art comparable novel-view synthesis results. Additionally, our neural-free sparse voxels are seamlessly compatible with grid-based 3D processing algorithms. We achieve promising mesh reconstruction accuracy by integrating TSDF-Fusion and Marching Cubes into our sparse grid system.
Figure 1: We propose SVRaster, a novel framework for multi-view reconstruction and novel view synthesis. (a) Sparse voxel representation effectively captures the volume density and radiance field of the scene, without the need for neural networks or sparse-points prior initialization. (b) Using our customized sparse voxel rasterizer, we can learn the underlying 3D scene efficiently and achieve state-of-the-art performance in both rendering quality and speed. (c) Notably, our sparse voxel representation can be naturally integrated with classic 3D algorithms such as TSDF-Fusion [7, 31, 8] and Marching Cubes [23], making it flexible and suitable for a wide range of applications.
@@ 1 Introduction
With these observations in mind, we present SVRaster, a novel framework that combines the efficiency of rasterization in 3DGS with the structured volumetric approach of grid-based representations. SVRaster leverages (1) multi-level sparse voxels to model 3D scenes and (2) implements a dynamic Morton order encoding that facilitates the rasterization rendering from our adaptive-sized sparse voxel representation. Our rendering is free from popping artifacts because the 3D space is partitioned into disjoint voxels, and our sorting ensures the correct rendering order. Moreover, thanks to the volumetric nature and neural-free representation of SVRaster, our sparse voxel grid can be easily and seamlessly integrated with classical grid-based 3D processing algorithms, as demonstrated with the incorporated sparse-voxel TSDF Fusion and Marching Cubes.
We show that our method is training fast, rendering fast, and achieves novel-view synthesis quality comparable to the state-of-the-art. We extract meshes by the integrated sparse-voxel TSDF-Fusion and sparse-voxel Marching Cubes from the trained sparse voxel grids. Results on mesh reconstruction benchmarks show our good balance between surface accuracy and processing speed.
@@ 2 Related Work
Differentiable volume rendering has made significant strides in 3D scene reconstruction and novel-view synthesis tasks. Neural Radiance Fields (NeRF) [27] laid the foundation for this progress by optimizing a volumetric function, parameterized by multi-layer perceptrons (MLPs) to encode both geometry and appearance through differential rendering. Subsequent studies focus on accelerating speed by decomposing large MLPs into grid-based representations, with a shallow MLP typically still being employed. Several grid representations have been explored, including dense grids [37], sparse grids [11], factorized grids [5], tri-planes [4], and hash grids [29]. The most similar scene representation to ours is the sparse grid by Plenoxels [11]. We both explicitly model 3D scenes into voxels without neural networks. Our quality and scalability are significantly improved due to adaptive and non-uniform voxels, while our rendering is much faster because of rasterization. Another scalability issue of the previous grid-based methods is that they still use some sort of dense 3D grid. For instance, dense occupancy grids [37, 5, 29, 21] for free-space skipping or dense pointer grid [11] to support sparse voxel lookup. We do not use any 3D dense data structures.
@@
Mesh reconstruction is an important extended topic for both volumetric-based and GS-based methods. Recent 3DGS variants [16, 14] with proper regularization have shown a good speed-accuracy trade-off. However, directly extracting meshes from 3DGS remains challenging, so these methods still rely on volumetric-based post-processing algorithms to extract meshes. NeRF ariants [41, 46, 38, 22, 45, 42] improve surface quality by rendering a signed distance field (SDF) instead of density field. The well-defined volume makes it straightforward to extract a mesh from the isosurface. Our sparse voxels are also a volumetric representation. Though our voxel also follows volumetric representation focusing on the density field, we still can achieve promising accuracy with minimal time.
For efficient rendering, another line of research is to convert (or distill) the geometry and appearance of a high-quality but slow model into other efficient kind of representations such as smaller MLPs [33, 40], meshes [43, 47, 34], grids [10], or the recent Gaussians [30]. Typically, the powerful Zip-NeRF [3] is employed as the teacher model which requires hours of training per scene. Note that using strong prior can be regarded as a different setup from ours.
Finally, the sparse voxel is also a popular representation for 3D processing and understanding [6, 13, 39, 44]. We mainly focus on fast and high-quality rendering while scene understanding is out of the scope of our work.
@@ 3.2.4 Sparse-voxel TSDF Fusion and Marching Cubes
Ours-small 240 8m 0.213 26.78 0.796
Ours-base 137 12m 0.197 27.10 0.807
Ours-large 92 17m 0.192 27.38 0.814
Re-evaluate on our machine using the public code.
Table 1: Novel-view synthesis results comparison on Mip-NeRF360 dataset [2]. The results are averaged from 4 indoor scenes and 5 outdoor scenes.
3DGS variants Ours
Method 3DGS [19] StopThePop [32] 3DGRT [28] EVER [24] base large
Rasterization
No popping $\triangle$ $\triangle$
[…]
† Re-evaluated on our machine using the public code.
Table 2: Comparison with recent 3DGS variants on Mip-NeRF360 dataset [2]. The LPIPS values here are evaluated with the correct intensity scale between $[-1,1]$ following EVER [24]. 3DGRT and EVER use ray tracing approach instead of rasterization, where only the concurrent EVER can totally resolve the popping artifact of GS-based rendering. Please find supplementary for more visual comparison.
@@
Ours 110 0.165 22.73 260 0.243 29.57
Re-evaluate on our machine using the public code.
@@ 4.1 Implementation Details
We use the following implementation details except stated otherwise. We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels (Secs. 3.2.1 and 5). The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.1$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for mesh reconstruction task. We train our model for 20K iterations. We find the merit of training longer is marginal to us. The voxels are pruned and subdivided every $h_{\mathrm{every}}{=}1{,}000$ iterations until 15K iterations. We prune voxels with less than $h_{\mathrm{prune}}{=}0.05$ maximum blending weights and subdivide top $h_{\mathrm{percent}}{=}5$ percent of voxels with higher priority (Eq. 8). More details are provided in the supplement.
@@ 4.2 Novel-view Synthesis
We show results on the real-world datasets in this section. Results on synthetic data are in supplementary materials.
In Tab. 1, we show the quantitative comparison on the MipNeRF-360 dataset. We mainly compare the representative scene representations with efficient training and rendering. Regarding frame rates, 3DGS [19] and ours with rasterization achieve much higher frame rates compared to the previous ray casting and sampling approach [11, 29]. Zip-NeRF [3] still achieves the best novel-view quality but is prohibitively slow to train and render. Compared to the competitive 3DGS [19], our base setup achieves similar FPS, better LPIPS, and training time, but worse in PSNR and SSIM. This may be explained by the qualitative comparison in Fig. 6. We find 3DGS rendering is typically smoother than ours due to its Gaussian representation while 3DGS sometimes fails to recover finer details. Some other important factors include: i) 3DGS uses the SfM sparse points to initialize while we do not and ii) we use different strategies to progressively grow our model. However, it is difficult to align everything for a side-by-side comparison due to the difference in our scene representations.
We also train a small variant by pruning more voxels with $h_{\mathrm{prune}}{=}0.15$ and a large variant by pruning less $h_{\mathrm{prune}}{=}0.01$, higher supersample rate $h_{\mathrm{ss}}{=}1.5$, and $K{=}2$ sample points inside each voxel. Our small variant achieves similar LPIPS with +80% FPS compared to 3DGS. Our large variant has 92 FPS with LPIPS close to Zip-NeRF.
More comparisons with recent 3DGS variants are in Tab. 2. 3DGS [19] has popping artifacts due to the ordering and overlapping issues. StopThePop [32] uses running sort and 3DGRT [28] uses ray tracing for accurate ordering but drops FPS by 28% and 67% respectively. EVER [24] further handles the Gaussian overlapping cases but with even less FPS. Our method ensures correct ordering (Sec. 3.1.2) with FPS and quality on par with the original 3DGS.
We also compare results on two more datasets in Tab. 3. On the Tanks&Temples dataset [20], despite having better LPIPS, our PSNR, and FPS are much worse than 3DGS. The failure results are shown in Fig. 8. We notice that the training views of the Tanks&Temples dataset have large exposure variation. As a result, our method produces a clear brightness boundary in the scene with many floaters that slow down our rendering. On the Deep Blending [15] dataset, we achieve similar LPIPS and PSNR as 3DGS but with 85% higher FPS. The speed and quality of 3DGS and our method are scene-dependent. More results and discussion are in the supplementary.
@@ 4.3 Ablation Studies
In our rasterizer, we use dynamic Morton order to ensure the sorting of voxels from different levels is correct. However, like the case in 3DGS, the popping artifact by the incorrect order is not a major factor in the numerical results. Instead, we provide more visualization in the supplementary videos to show its effect. See more ablation studies in the supplementary materials.
@@ 4.4 Mesh Reconstruction
We use sparse-voxel TSDF-Fusion and Marching Cubes (Sec. 3.2.4) to extract mesh from the optimized volume density and sparse grid layout. The results comparison on the large-scale Tanks&Temples and the object-scale DTU datasets is provided in Tab. 5. Neuralangelo [22] is still state-of-the-art in terms of mesh accuracy but needs hours of optimization per scene. Our method achieves a good balance of accuracy and training time on both datasets. Interestingly, our method with density modeling still achieves on-par geometric accuracy with some other methods using a signed distance field (SDF). This shows that a density field with proper regularization and strong scene modeling capacity could still potentially achieve good accuracy on both large-scale and object-scale scenes. Compared to the recent GS-based methods, we do not use the coarse geometry estimated by SfM and still achieve better accuracy and training time. Some qualitative examples are in Fig. 7. More results are in the supplementary materials.
@@ 5 Conclusion
This work presents a novel differentiable radiance field rendering system that integrates an efficient rasterizer with a multi-level sparse voxel scene representation. We reconstruct a scene from the input multi-view images by adaptively fitting the sparse voxels into different levels-of-detail. We also incorporate some classical grid-based 3D processing algorithms into our system including sparse-voxel TSDF-Fusion and Marching Cubes for mesh extraction. Experiments on novel-view synthesis and mesh reconstruction tasks demonstrate our effectiveness, achieving high FPS rendering, fast training, and high result quality.
@@ B More Details of Voxel Rendering Order
Our sorting-based rasterizer is based on the efficient CUDA implementation done by 3DGS [19]. In the following, we first describe how the overall sorting pipeline works in Sec. B.1. We then dive more into the implementation of the dynamic Morton order in Sec. B.2 and its correctness proof in Sec. B.3.
@@ B.1 Overview
In our case, we replace the primitive z-depth with a dynamic Morton order of voxels to ensure the rendering order is always correct. As there are eight different Morton orders to follow depending on the positive/negative signs of ray directions, dubbed ray sign bits, we further duplicate each voxel by the numbers of different ray sign bits it covers. The ray sign bits are also attached to each duplicated voxel. In the rendering stage, a pixel only composites voxels with the same attached ray sign bits when there are multiple ray sign bits in an image tile. Our bit field of the key-value pair is:
key $=|\underbrace{\text{tile id}}_{\text{16 bits}}|\underbrace{\text{dynamic Morton order}}_{\text{48 (=$3L$) bits}}|$ (24a)
@@
def to_dynamic_morton_order(octpath, signbits): # Input # octpath: Voxel Octree Morton code. # signbits: The signbits the voxel care. # Output # order: The order for sorting. table = order_tables[signbits] order = 0 for i in range(MAX_NUM_LEVELS): order —= table[octpath & 0b111] ¡¡ (3*i) octpath = octpath ¿¿ 3 return order Pseudocode for dynamic Morton order. The mapping between voxel grid $(i,j,k)$ index and Octree Morton code octpath is detailed in Sec. A.1. In practice, the mapping from Octree Morton code to dynamic Morton order is done by a single bitwise $\mathrm{xor}$ operation instead of for-loop. More details in Sec. B.2.
B.2 Dynamic Morton Order

### v3

@@ Abstract
We propose an efficient radiance field rendering algorithm that incorporates a rasterization process on adaptive sparse voxels without neural networks or 3D Gaussians. There are two key contributions coupled with the proposed system. The first is to adaptively and explicitly allocate sparse voxels to different levels of detail within scenes, faithfully reproducing scene details with $65536^{3}$ grid resolution while achieving high rendering frame rates. Second, we customize a rasterizer for efficient adaptive sparse voxels rendering. We render voxels in the correct depth order by using ray direction-dependent Morton ordering, which avoids the well-known popping artifact found in Gaussian splatting. Our method improves the previous neural-free voxel model by over 4db PSNR and more than 10x FPS speedup, achieving state-of-the-art comparable novel-view synthesis results. Additionally, our voxel representation is seamlessly compatible with grid-based 3D processing techniques such as Volume Fusion, Voxel Pooling, and Marching Cubes, enabling a wide range of future extensions and applications. Code: github.com/NVlabs/svraster
Figure 1: We propose SVRaster, a novel framework for multi-view reconstruction and novel view synthesis. (a) Sparse voxel representation effectively captures the volume density and radiance field of the scene, without the need for neural networks, 3D Gaussians, and sparse-points prior. (b) Using our customized sparse voxel rasterizer, we can learn the underlying 3D scene efficiently and achieve state-of-the-art performance in both rendering quality and speed. (c) Notably, lifting 2D modal to the trained sparse voxels is simple and efficient by integrating the classic Volume Fusion [7, 36, 8]. We show examples of vision foundation model feature field from RADIO [40], semantic field from Segformer [55], and signed distance field from rendered depth, making it flexible and suitable for a wide range of applications.
@@ 1 Introduction
With these observations in mind, we present SVRaster, a novel framework that combines the efficiency of rasterization in 3DGS with the structured volumetric approach of grid-based representations. SVRaster leverages (1) multi-level sparse voxels to model 3D scenes and (2) implements a direction-dependent Morton order encoding that facilitates the rasterization rendering from our adaptive-sized sparse voxel representation. Our rendering is free from popping artifacts because the 3D space is partitioned into disjoint voxels, and our sorting ensures the correct rendering order. Moreover, thanks to the volumetric nature and neural-free representation of SVRaster, our sparse voxel grid can be easily and seamlessly integrated with classical grid-based 3D processing algorithms.
We show that our method is training fast, rendering fast, and achieves novel-view synthesis quality comparable to the state-of-the-art. We also integrate Volume Fusion, Voxel Pooling, and Marching Cubes operations into our adaptive sparse voxels, which showcases promising results of mesh extraction and 2D foundation feature fusion.
@@ 2 Related Work
Differentiable volume rendering has made significant strides in 3D scene reconstruction and novel-view synthesis tasks. Neural Radiance Fields (NeRF) [32] laid the foundation for this progress by optimizing a volumetric function, parameterized by multi-layer perceptrons (MLPs) to encode both geometry and appearance through differential rendering. Subsequent studies focus on accelerating speed by decomposing large MLPs into grid-based representations, with a shallow MLP typically still being employed. Several grid representations have been explored—dense grids [45], factorized grids [5], tri-planes [4], and hash grids [34].
In particular, our method is closely related to sparse voxel grid representations. Sparse Voxel Octrees [22] and VDB trees [35] are commonly used to manage sparse voxels and facilitate rendering. In contrast, our sparse voxels are stored in a 1D array without advanced data structures. Our rasterizer with the proposed direction-dependent Morton order ensures correct rendering order. To model volumes in sparse leaf nodes, different strategies exist, such as low-resolution 3D grid [35, 53] or implicit neural field [25, 47]. Our leaf node is a single voxel with explicit density and color parameters, like Plenoxels [11]. Regarding voxel levels, previous methods allocate voxels to a target level [25, 47, 59, 11] or use a shallow tree [35, 53]. Our voxels can adaptively fit into different levels across the entire tree depth, maximizing flexibility and scalability.
Another scalability issue of the previous grid-based methods is that they still use some sort of dense 3D grid in the field of novel-view synthesis. For instance, dense occupancy grids [45, 5, 34, 23] for free-space skipping or dense pointer grid [11] to support sparse voxel lookup. We do not use any 3D dense data structures.
@@
Mesh reconstruction is an important extended topic for both volumetric-based and GS-based methods. Recent 3DGS variants [16, 14] with proper regularization have shown a good speed-accuracy trade-off. However, directly extracting meshes from 3DGS remains challenging, so these methods still rely on volumetric-based post-processing algorithms to extract meshes. NeRF variants [50, 56, 46, 24, 54, 51] improve surface quality by rendering a signed distance field (SDF) instead of density field. The well-defined volume makes it straightforward to extract a mesh from the isosurface. Our sparse voxels are also a volumetric representation. Though our voxel also follows volumetric representation focusing on the density field, we still can achieve promising accuracy with minimal time.
For efficient rendering, another line of research is to convert (or distill) the geometry and appearance of a high-quality but slow model into other efficient kind of representations such as smaller MLPs [41, 49], meshes [52, 57, 42], grids [10], or the recent Gaussians [37]. Typically, the powerful Zip-NeRF [3] is employed as the teacher model which requires hours of training per scene. Our volumetric representation could enable training-free conversion from a large implicit field, which is however out of our current scope.
Finally, the sparse voxel is also a widely-used representation for 3D processing [19, 7, 36, 8, 38, 26] and understanding [6, 13, 53]. We mainly focus on rendering in this work while we believe that adapting our method to these techniques is a promising future direction.
@@ 3.2.4 Sparse-voxel TSDF Fusion and Marching Cubes
Ours fast-rend 258 9m 0.210 26.87 0.804
Ours fast-train 131 4.5m 0.199 27.08 0.816
Ours 121 15m 0.185 27.33 0.822
Re-evaluate on our machine.
Table 1: Novel-view synthesis results comparison on Mip-NeRF360 dataset [2]. The results are averaged from 4 indoor scenes and 5 outdoor scenes. 3DGS uses the sparse points prior from COLMAP [43], whereas the other methods and ours do not.
@@
Ours 125 0.144 23.04 366 0.228 29.84
Re-evaluate on our machine.
@@ 4.1 Implementation Details
We use the following implementation details except stated otherwise. We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels (Secs. 3.2.1 and 5). The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.5$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for mesh reconstruction task. We train our model for 20K iterations. The voxels are pruned and subdivided every $h_{\mathrm{every}}{=}1{,}000$ iterations. The pruning threshold $h_{\mathrm{prune}}$ is linearly scale from $0.0001$ to $0.05$ and the subdivision percentage is $h_{\mathrm{percent}}{=}5$. More details are provided in the supplement.
@@
FPS at higher res.$\uparrow$
Method Peak GPU mem.$\downarrow$ Model $\downarrow$ 1x 2x 3x
3DGS 1.8GB 0.7GB 131 69 39
Ours 3.9GB 1.8GB 121 103 69
Table 3: Memory, model size, and high-res FPS. The results are averaged on Mip-NeRF360 dataset. The standard 1x evaluation resolution have about 1–1.6M pixels per frame.
@@ 4.2 Novel-view Synthesis
In Tab. 1, we show the quantitative comparison on the MipNeRF-360 dataset. We also provide a fast-rendering variant with >2x FPS by setting $h_{\mathrm{ss}}{=}1.5,h_{\mathrm{prune}}{=}0.15$, and a fast-training variant with >3x training speedup by scaling all iteration related hyperparameter by $0.3$. Both fast variants only reduce quality moderately.
Our rendering speed is comparable to 3DGS on average. However, FPS varies significantly across scenes (see supplementary for detailed per-scene results). Several differences impact speed between our method and 3DGS. For instance, 3DGS sorts 32-bit floats, while we sort 48-bit Morton codes (16 levels, each with 3 bits) for primitive ordering. We also avoid the overhead of computing inverse covariance matrices from quaternions and scaling parameters. We directly intersect rays with voxels in 3D space, whereas 3DGS approximates the projection of 3D Gaussians to 2D via a linear affine transform. Finally, Gaussian distributions decay gradually to zero, while our post-activation [45, 18] voxel density field can be arbitrarily sharp. As a result, the average numbers of primitives contributing to a pixel intensity is 63 for 3DGS and 27 for us. The influence of these factors on rendering speed is scene-dependent.
Regarding rendering quality metrics, our method achieves significantly better LPIPS than 3DGS and even surpasses Zip-NeRF. This can be explained by the qualitative comparison in Fig. 6 where our method recover finer details. The SSIM comparison results with 3DGS is more scene-dependent. SSIM comparisons with 3DGS are more scene-dependent, while PSNR tends to favor 3DGS, as it prefers smoother rendering in uncertain regions. More visual comparisons are in the supplementary.
Our method uses much more voxel primitives than Gaussian, resulting in larger model size and requires more GPU memory as shown in Tab. 3. Interestingly, our high FPS is more scalable to higher resolution rendering perhaps due to the fewer contributing primitives per pixel.
We compare results on two more datasets in Tab. 2. On Deep Blending [15] dataset, we achieve much better speed and quality than 3DGS. On Tanks&Temples dataset [21], despite having better LPIPS, our PSNR and FPS are worse than 3DGS. The failure results are shown in Fig. 8 where the training views have large exposure variation. As a result, our method produces a clear brightness boundary in the scene with many floaters that slow down our rendering.
@@ 4.3 Ablation Studies
In our rasterizer, we use direction-dependent Morton order to ensure the sorting of voxels from different levels is correct. However, like the case in 3DGS, the popping artifact by the incorrect order is not a major factor in the numerical results. Instead, we provide more visualization in the supplementary videos to show its effect. See more ablation studies in the supplementary materials.
@@ 4.4 Mesh Reconstruction
We extract mesh by adapting TSDF-Fusion [36] and Marching Cubes [27] into our sparse voxels. The results comparison on the large-scale Tanks&Temples and the object-scale DTU datasets is provided in Tab. 5, where our method achieves a good balance of accuracy and training time on both datasets. See qualitative examples in Fig. 7 and the supplementary materials. We find that our current method tends to produce unnecessary geometric bumps for texture details. Future work could improve surface smoothness by adapting our density field for direct signed distance field modeling or incorporating SfM sparse points prior.
4.5 2D Feature Fusion
We showcase some results in Fig. 1, where we fuse 2D semantic segmentation and high-dimensional foundation features into our sparse voxels using Voxel Pooling and Volume Fusion. The multi-view ensemble smooths the inconsistent 2D predictions, while our detailed 3D geometry enables higher-resolution rendering of 2D features. See more results in the released code.
@@ 5 Conclusion
This work presents a novel differentiable radiance field rendering system that integrates an efficient rasterizer with a multi-level sparse voxel scene representation. We reconstruct a scene from the multi-view input images by adaptively fitting the sparse voxels into different levels-of-detail. The results reveal that fully explicit voxel models, without neural networks or Gaussians, can achieve state-of-the-art comparable novel-view rendering speed and quality. The key breakthrough is overcoming the scalability constraint of voxel grids with adaptive sparse voxels and improving rendering speed through rasterization. We believe that incorporating our method with classical 3D processing algorithm [19, 7, 8, 38, 26] and 3D neural network [6, 13, 53] are promising future directions.
Acknowledgements.
We thank Min-Hung Chen and Yueh-Hua Wu for helpful paper proofreading.
Updates.
  • Mar 2025: (1) Revise literature review. (2) Add Scannet++ results.
  • Feb 2025: (1) Code release. (2) Add more volume fusion examples. (3) Novel-view synthesis quality improved. (4) Provide fast rendering and fast training variants. (5) Discuss results more.
  • Dec 2024: Paper release on arxiv.
@@ B More Details of Voxel Rendering Order
Our sorting-based rasterizer is based on the efficient CUDA implementation done by 3DGS [20]. In the following, we first describe how the overall sorting pipeline works in Sec. B.1. We then dive more into the implementation of the direction-dependent Morton order in Sec. B.2 and its correctness proof in Sec. B.3.
@@ B.1 Overview
In our case, we replace the primitive z-depth with a direction-dependent Morton order of voxels to ensure the rendering order is always correct. As there are eight different Morton orders to follow depending on the positive/negative signs of ray directions, dubbed ray sign bits, we further duplicate each voxel by the numbers of different ray sign bits it covers. The ray sign bits are also attached to each duplicated voxel. In the rendering stage, a pixel only composites voxels with the same attached ray sign bits when there are multiple ray sign bits in an image tile. Our bit field of the key-value pair is:
key $=|\underbrace{\text{tile id}}_{\text{16 bits}}|\underbrace{\text{Morton order}}_{\text{48 (=$3L$) bits}}|$ (24a)
@@
def to_dir_dep_morton_order(octpath, signbits): # Input # octpath: Voxel Octree Morton code. # signbits: The signbits the voxel care. # Output # order: The order for sorting. table = order_tables[signbits] order = 0 for i in range(MAX_NUM_LEVELS): order —= table[octpath & 0b111] ¡¡ (3*i) octpath = octpath ¿¿ 3 return order Pseudocode for direction-dependent Morton order. The mapping between voxel grid $(i,j,k)$ index and Octree Morton code octpath is detailed in Sec. A.1. In practice, the mapping from Octree Morton code to direction-dependent Morton order is done by a single bitwise $\mathrm{xor}$ operation instead of for-loop. More details in Sec. B.2.
B.2 Direction-dependent Morton Order